OpenStudy (anonymous):

A powerful missile reaches a speed of 5 kilometers per second in 10 seconds after its launch. What is the average acceleration of the missile during this period? a) 0.2 meters/second^2 b) 0.4 centimeters/second^2 c) 0.5 meters/second^2 d) 0.5 kilometers/second^2

OpenStudy (shane_b):

Use:\[\large\frac{\Delta V}{\Delta t}=\frac{V_f-V_i}{10s}=\frac{5km/s-0km/s}{10s}=0.5km/s^2\]

OpenStudy (anonymous):

I got that answer by guessing but I don't understand what the V are for.

OpenStudy (shane_b):

velocity (final) and velocify(initial)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (shane_b):

np :) Good luck